A school district is trying to decrease the uncertainty concerning pick-up times on a particular school bus route. Due to uncertainty in traffic patterns, the school district examines two alternative routes. In a sample of 15 days on alternative route #1, the average driving time for the school bus is 40 minutes with a standard deviation of 15 minutes. In a sample of 16 days using alternative route #2, the average drive time is 44 minutes with a standard deviation of 11 minutes.

At the 5% level of significance, is there a difference between the driving times of the two bus routes. Show all work on how you come to your conclusion and provide a clear explanation as if you were explaining the results to the school district.
Answer #1

Solution

The solution is based on two-sample t-test.

Let X = pick-up time (minutes)in route #1

      Y = pick-up time (minutes)in route #2

Let mean and standard deviation of X be respectively µ1 and σ1 and those of Y be µ2 andσ2, where σ12 = σ22 = σ2, say and σ2 is unknown.

Hypotheses:

Null: H0: µ1 = µ2 Vs Alternative: HA: µ1 µ2

Test Statistic:

t = (Xbar - Ybar)/[s√{(1/n1) + (1/n2)}]

where

s2 = {(n1 – 1)s12 + (n2 – 1)s22}/(n1 + n2 – 2);

Xbar and Ybar are sample averages and

s1,s2 are sample standard deviations based on n1 observations on X and n2 observations on Y respectively.

Calculations

Summary of Excel calculations is given below:

n1

15

n2

16

Xbar

40

Ybar

44

s1

15

s2

11

s2

171.2069

s

13.08461

tcal

0.850598

α

0.05

tcrit

2.04523

p-value

0.401962

(n1 + n2 - 2)

29

0.1

n1 -1

14

n2 -1

15

s12

225

s22

121

(n1 - 1)s12

3150

(n2 - 1)s22

1815

sum

4965

n1 + n2-2

29

1/n1

0.066667

1/n2

0.0625

sum

0.129167

sqrt

0.359398

s.sqrt (SE)

4.702576

Distribution, Significance Level, α Critical Value and p-value:

Under H0, t ~ tn1 + n2 - 2. Hence, for level of significance α%,

Critical Value = upper (α/2)% point of tn1 + n2 - 2 and

p-value = P(tn1 + n2 - 2 > | tcal |)

Using Excel Function: Statistical TINV andTDIST, these are found to be as shown in the above table.

Decision:

Since | tcal | < tcrit, or equivalently since p-value > α, H0 is accepted.

Conclusion:

There is not sufficient evidence to suggest that the mean pick-up times differ.

Hence, conclude that

there is no difference between the driving times of the two bus routes. Answer 1

DONE